A well lagged aluminium calorimeter of mass 80g contains 150g of water and 100g of ice at 0^oC. A heating coil rated 1.0kw is put in the calorimeter and the mixture is stirred until its temperature is 33^oC. Calculate how much ice is left after 1 minute. State any assumption you made.
1
Expert's answer
2021-02-22T10:29:54-0500

Mass of aluminum, "m_a = 80g = 0.08Kg"

Mass of water, "m_w = 150 g = 0.15 Kg"

Mass of ice, "m_i = 100g = 0.1 Kg"

Specific heat of water, "s_w = 4182K\/Kg\/K"

Specific heat of aluminum, "s_a = 0.9 J\/g K = 900J\/Kg\/K"

Specific heat of ice, "s_i = 2.04KJ\/Kg\/K = 2040J\/Kg\/K"

Latent heat of ice, "L_f = 3.33 \times 10^5 J/Kg"

Initial temperature, "T_i = 0^\circ C"

Final temperature, "T_f = 33^\circ C"


Heat generated by the coil in one minute, "Q = 1000 \times 60 J = 60000 J"

Let x amount of the ice that melt,

then assume that there is no loss of the heat,

"(m_a s_a+m_ws_w+xs_i)(T_f-T_i)+m_iL_f =60000"

Putting values,

"[(0.080 \times 900)+(0.15 \times 4182)+(x \times 2040)](33-0) +(0.1 \times 3.33 \times 10^5) = 60000"


Solving it, we get, "x =0.05 Kg = 50 g"


Amount of ice left, "M = 100-50 = 50 g"